OpenStudy (anonymous):

Solve the following system of equations. x - 2y = 5 2x - 4y = 10 Infinitely Many Solutions No Solutions (0, 0) (5, 10)

hero (hero):

x + 5 = 2y 2x + 10 = 4y y = (x+5)/2 y = (2x+10)/4 4(x+5) = 2(2x+10) 4x + 20 = 4x + 20 Infinite Solutions*